# Achieving Near-Optimal Regret for Bandit Algorithms with Uniform Last-Iterate Guarantee

### v1

@@ Appendix G Omitted Details of Section 4
Before proving our main result of Algorithm 3, we first prove some technical lemmas in the following subsections.
G.1 Proof of Lemma 4.6
Lemma G.1 (Restatement of Lemma 4.6).
@@
Moreover, according to Lemma 4.6, for all $m\geq 2$, $S_{m}$ can also be equivalently rewritten as:
@@
where the first equality holds due to Lemma 4.6 and the last equality holds due to Eq. (47).
Step 2: ${\mathcal{B}}_{m}^{\prime}$ is a $C$-approximate barycentric spanner of ${\mathcal{S}}_{m}^{\prime}$. The proof idea of this step follows a similar arguments of Proposition 2.2 and Proposition 2.5 in (Awerbuch and Kleinberg, 2008). In Algorithm 4, each update for matrix $A\in\mathbb{R}^{d\times d}$ always maintains $|\det(A)|>0$, and thus ${\mathcal{B}}_{m}^{\prime}$ spans $\mathbb{R}^{d}$. As ${\mathcal{S}}_{m}^{\prime}\subseteq\mathbb{R}^{d}$, every element $a\in{\mathcal{S}}_{m}^{\prime}$ can be linearly represented by $a=\sum_{i=1}^{d}w_{i}b_{i}$ for some coefficients $\{w_{i}\}_{i=1}^{d}$, and then we have
@@
G.3 Proof of Theorem 4.2: Regret Analysis
@@
[…]
@@
[…]
@@
Notice that the above inequality holds for all $m\in\mathbb{N}$. Finally, Theorem 20.5 of (Lattimore and Szepesvári, 2020) shows that with probability at least $1-\delta$, for all $m\in\mathbb{N}$
@@
G.4 Proof of Theorem 4.2: Computational Analysis
@@
Under the same setting of Lemma 4.3, Algorithm 4 outputs a $C$-barycentric spanner by restarting the second for-loop for $\mathcal{O}\left(d^{2}\log_{C}(d)\right)$ times.
@@
According to (Awerbuch and Kleinberg, 2008, Lemma 2.6), if ${\mathcal{I}}_{m}$ never changes after entering the second for-loop, then, the second for-loop restarts for at most $\mathcal{O}(d\log_{C}(d))$ times, and then the algorithm terminates. In Algorithm 4, if ${\mathcal{I}}_{m}$ changes, the second for-loop restarts. As set ${\mathcal{I}}_{m}$ is always non-increasing, it suffices to consider the worst case (${\mathcal{I}}_{m}$ changes at most $O(d)$ times). Hence, the second for-loop restarts at most $O(d^{2}\log_{C}(d))$ times. ∎

### v2

@@ Appendix G Omitted Details of Section 4
G.1 Notations
We use $e_{i}\in\mathbb{R}^{d}$ to denote a vector whose $i$-th coordinate is one and all others are zero. For any two vectors $x,y\in\mathbb{R}^{d}$, $x\leq y$ indicates that $x_{i}\leq y_{i}$ holds for each coordinate $i$. For a positive definite matrix $A\in\mathbb{R}^{d\times d}$, the weighted $2$-norm of vector $x\in\mathbb{R}^{d}$ is given by $\left\|{x}\right\|_{A}=\sqrt{x^{\top}Ax}$. For matrix $A=[a_{1},\cdots,a_{d}]\in\mathbb{R}^{d\times d}$ with each $a_{i}\in\mathbb{R}^{d}$, we use $A_{-i}$ to denote the $(d-1)$-tuple of vectors $[a_{1},\cdots,a_{i-1},a_{i+1},\cdot,a_{d}]$. For matrices $A,B$, we use $A\succ B$ to indicate that $A-B$ is positive definite. For two sets ${\mathcal{A}},{\mathcal{B}}$, we use ${\mathcal{A}}-{\mathcal{B}}$ to indicate the exclusion. For any scalar $C\in\mathbb{R}$, we use $\vec{C}$ to denote a vector, with all coordinates equal to $C$.
G.2 Key Technique: Adaptive Barycentric Spanner
Algorithm 12 Adaptive barycentric spanner
if $|{\mathcal{B}}_{m-1}|<d$ then
      3 Use Gaussian elimination to get $M_{m}\in\mathbb{R}^{(d-|{\mathcal{B}}_{m-1}|)\times d}$ such that $\text{Span}({\mathcal{B}}_{m-1})=\{x\in\mathbb{R}^{d}:M_{m}x=\vec{0}\}$.
4else
      5 Remove constraints in (42), (13) related to $B_{m-1},M_{m}$.
76 Query oracle to get empirical best arm $a^{\star}_{m}$, the solution of
$\quad\operatorname*{\arg\!\max}_{a\in{\mathcal{A}}}\left\langle{\widehat{\theta}_{m},a}\right\rangle$ (42)
@@ 76 Query oracle to get empirical best arm $a^{\star}_{m}$, the solution of
for $i=1,\ldots,d$ do
      98 Set $s_{i}=\text{LI-Argmax}({\mathcal{A}},A,\widehat{\theta}_{m},{\mathcal{B}}_{m-1},m,C,a^{\star}_{m})$. if $s_{i}\neq\textsc{Null}$ then
            10 Update $a_{i}=s_{i}$ and ${\mathcal{I}}_{m}={\mathcal{I}}_{m}-\{i\}$.
11 for $i=1,\ldots,d$ do
      1312 Set $s_{i}=\text{LI-Argmax}({\mathcal{A}},A,\widehat{\theta}_{m},{\mathcal{B}}_{m-1},m,C,a^{\star}_{m})$. if $s_{i}\neq\textsc{Null}$ then
            14 if $\left|\det\left(s_{i},A_{-i}\right)\right|\geq C\left|\det(A)\right|$ $\mathbf{or}$ $i\in{\mathcal{I}}_{m}$ then
                  1615 Update ${\mathcal{I}}_{m}={\mathcal{I}}_{m}-\{i\}$ and $a_{i}=s_{i}$. Restart this for-loop with current parameters.
17 Return: $\{a_{i}\}_{i=1}^{d}-\cup_{i\in{\mathcal{I}}_{m}}\frac{e_{i}}{\sqrt{T_{m}}}$.
Algorithm 12 initializes a matrix $A=[a_{1},\ldots,a_{d}]$ whose column vectors are linearly independent, and $a_{i}=e_{i}/\sqrt{T_{m}}$. At the beginning, the algorithm solves (42) to find an empirical best arm $a^{\star}_{m}$ (line 12-12). Then, in two for-loops (line 12-12 and line 12-12), the algorithm tries to replace each column vector of $A$ with an active arm $a\in{\mathcal{A}}_{m}$ by invoking the subroutine LI-Argmax in Algorithm 13, while keeping column vectors of $A$ linearly independent. For each column $i$, LI-Argmax attempts to find $\max_{a\in{\mathcal{A}}_{m}}|\det(a,A_{-i})|$ [^8] It is equivalent to $\max_{a\in{\mathcal{A}}_{m}}|\left\langle{u,a}\right\rangle|$ where $u$ can be found by rank-one update. See [Zhu et al., 2022] for more details. by solving (13) with $a^{\star}_{m}$. If $\max_{a\in{\mathcal{A}}_{m}}|\det(a,A_{-i})|=0$, then, the algorithm returns Null, since there does not exist an active arm that can maintain the linear independence. In this case, the algorithm keeps $e_{i}/\sqrt{T_{m}}$ in the $i$-th column of $A$. Otherwise, $a_{i}$ will be updated to $s_{i}\in\operatorname*{\arg\!\max}_{a\in{\mathcal{A}}_{m}}|\det(a,A_{-i})|$. We use set ${\mathcal{I}}_{m}$ to record the indices of columns where the replacement fails. According to the conditions in line 12, the second for-loop will be repeatedly restarted to ensure that the linear combination of elements in ${\mathcal{B}}_{m}$ can represent all active arms in ${\mathcal{A}}_{m}$ with coefficients in the range of $[-C,C]$. Also note that, as shown in Theorem 4.2, LI-Argmax will be oracle-efficient as it only queries the oracle $\widetilde{\mathcal{O}}\left(d^{3}\right)$ times.
Formally, we have the following result.
Lemma G.1.
Algorithm 13 Linear-independent argmax (LI-Argmax)
Query oracle to obtain $a^{+}$, the solution of
$\quad\operatorname*{\arg\!\max}_{a\in{\mathcal{A}}}\left\langle{u,a}\right\rangle$
s.t. $\left\langle{\widehat{\theta}_{m},a^{\star}_{m}-a}\right\rangle\leq 2^{-m+1},{}$ (43)
$M_{m}a=\vec{0},$ $\ -\vec{C}\leq\left(B_{m-1}^{\top}B_{m-1}\right)^{-1}B_{m-1}^{\top}a\leq\vec{C}.$
Query oracle to get $a^{-}$ from (13) by replacing $u$ with $-u$.
       Return: Null.
5else
       Return: $\operatorname*{\arg\!\max}_{b\in\left\{a^{+},a^{-}\right\}}\left|\left\langle{u,b}\right\rangle\right|$.
G.3 Computational Analysis
Definition G.2 (Optimization oracle).
We note that the constrained optimization oracles are commonly-used and also crucial for elimination-type approaches e.g., [Bibaut et al., 2020, Li et al., 2022]. Although this oracle is slightly powerful than that used for non-elimination based approaches [Dani et al., 2008, Agarwal et al., 2014], linear constrained oracle, in fact, can be implemented efficiently in many cases (e.g., a common assumption that ${\mathcal{A}}$ is a ball [Plevrakis and Hazan, 2020]) via the ellipsoid method. We refer readers to [Bibaut et al., 2020, Plevrakis and Hazan, 2020] for more discussions.
However, the evolution of phases blows up the number of constraints in (45) (e.g., when $m=\Omega(\exp(d))$), which prevents LI-Argmax from directly querying the optimization oracle defined in Definition G.2, since it requires the number of constraints at most $\mathcal{O}(d)$. To address this issue, one needs to remove the dependence on $m$.
Definition G.3 ($C$-bounded spanner).
For any given set ${\mathcal{S}}\subseteq\mathbb{R}^{d}$ and constant $C>0$, $\text{Span}_{[-C,C]}({\mathcal{S}})$, defined as follows, is a set that contains all possible linear combinations from ${\mathcal{S}}$ with coefficients within $[-C,C]$.
$\text{Span}_{[-C,C]}({\mathcal{S}})=\left\{\sum_{s\in{\mathcal{S}}}c_{s}\cdot s:\forall c_{s}\in[-C,C]\right\},$ (46)
where $c_{s}$ is the coefficient associated with element $s$.
We enlarge ${\mathcal{A}}_{m-1}$ to ${\mathcal{A}}\cap\text{Span}_{[-C,C]}({\mathcal{B}}_{m-1})$, and one can easily verify that this is true since ${\mathcal{A}}_{m-1}\cap{\mathcal{A}}={\mathcal{A}}_{m-1}$ and Lemma G.1 implies that ${\mathcal{A}}_{m-1}\subseteq\text{Span}_{[-C,C]}({\mathcal{B}}_{m-1})$.
Lemma G.4.
Suppose that $|{\mathcal{B}}_{m-1}|<d$. For $M_{m}$ and $B_{m-1}$ given in Algorithm 12, $\text{Span}_{[-C,C]}({\mathcal{B}}_{m-1})$ can be equivalently written as
Hence, LI-Argmax can invoke the optimization oracle in Definition G.2 to solve problems of $\operatorname*{\arg\!\max}_{a\in{\mathcal{A}}_{m-1}}\langle\widehat{\theta}_{m},a\rangle$ and (44) approximately by solving (42) and (13), respectively.
Computational efficiency. We mainly focus on two costly steps in Algorithm 12. The first step is to use Gaussian elimination to obtain the matrix $M_{m}\in\mathbb{R}^{n\times d}$ for some $n\leq d$. This step can be done with at most $\mathcal{O}(d^{3})$ complexity and is implemented at most once for each phase $m$. The other expensive step is to query the optimization oracle to construct a $C$-approximate barycentric spanner. The number of calls to the oracle depends on the number of times that Algorithm 12 invokes LI-Argmax subroutine. The LI-Argmax will be invoked for $d$ times in the first for-loop. For the second for-loop, LI-Argmax will be invoked for $\widetilde{\mathcal{O}}\left(d^{3}\right)$ times. Compared with $\widetilde{\mathcal{O}}\left(d^{2}\right)$ in [Awerbuch and Kleinberg, 2008], the extra $d$ in our complexity comes from the additional restart-condition $i\in{\mathcal{I}}_{m}$ in line 12 of Algorithm 12. As $|{\mathcal{I}}_{m}|\leq d$ and ${\mathcal{I}}_{m}$ is non-increasing when update, this condition will be met for at most $d$ times.
G.4 Proof of Lemma G.4
Lemma G.5 (Restatement of Lemma G.4).
Suppose that $|{\mathcal{B}}_{m-1}|<d$. For $M_{m}$ and $B_{m-1}$ given in Algorithm 12, the set $\text{Span}_{[-C,C]}({\mathcal{B}}_{m-1})={\mathcal{H}}_{m}$ where
@@ G.3 Computational Analysis
Moreover, according to Lemma G.4, for all $m\geq 2$, $S_{m}$ can also be equivalently rewritten as:
@@
where the first equality holds due to Lemma G.4 and the last equality holds due to Eq. (48).
Step 2: ${\mathcal{B}}_{m}^{\prime}$ is a $C$-approximate barycentric spanner of ${\mathcal{S}}_{m}^{\prime}$. The proof idea of this step follows a similar arguments of Proposition 2.2 and Proposition 2.5 in [Awerbuch and Kleinberg, 2008]. In Algorithm 12, each update for matrix $A\in\mathbb{R}^{d\times d}$ always maintains $|\det(A)|>0$, and thus ${\mathcal{B}}_{m}^{\prime}$ spans $\mathbb{R}^{d}$. As ${\mathcal{S}}_{m}^{\prime}\subseteq\mathbb{R}^{d}$, every element $a\in{\mathcal{S}}_{m}^{\prime}$ can be linearly represented by $a=\sum_{i=1}^{d}w_{i}b_{i}$ for some coefficients $\{w_{i}\}_{i=1}^{d}$, and then we have
@@
G.6 Proof of Theorem 4.2: Regret Analysis
@@
$\sup_{a\in{\mathcal{A}}}\left\|{a}\right\|^{2}_{V(\pi;{\mathcal{A}})^{-1}}\leq C\cdot d,\quad\text{where}\quad V(\pi;{\mathcal{A}})=\sum_{a\in{\mathcal{A}}}\pi(a)aa^{\top}.$
@@
$\sup_{a\in{\mathcal{A}}}\left\|{a}\right\|^{2}_{V(\pi;{\mathcal{B}})^{-1}}\leq C^{2}\cdot d^{2},\quad\text{where}\quad V(\pi;{\mathcal{B}})=\sum_{b\in{\mathcal{B}}}\pi(b)bb^{\top}.$
@@
Notice that the above inequality holds for all $m\in\mathbb{N}$. Finally, Theorem 20.5 of [Lattimore and Szepesvári, 2020] shows that with probability at least $1-\delta$, for all $m\in\mathbb{N}$
@@
G.7 Proof of Theorem 4.2: Computational Analysis
@@
Under the same setting of Lemma G.1, Algorithm 12 outputs a $C$-barycentric spanner by restarting the second for-loop for $\mathcal{O}\left(d^{2}\log_{C}(d)\right)$ times.
@@
According to [Awerbuch and Kleinberg, 2008, Lemma 2.6], if ${\mathcal{I}}_{m}$ never changes after entering the second for-loop, then, the second for-loop restarts for at most $\mathcal{O}(d\log_{C}(d))$ times, and then the algorithm terminates. In Algorithm 12, if ${\mathcal{I}}_{m}$ changes, the second for-loop restarts. As set ${\mathcal{I}}_{m}$ is always non-increasing, it suffices to consider the worst case (${\mathcal{I}}_{m}$ changes at most $O(d)$ times). Hence, the second for-loop restarts at most $O(d^{2}\log_{C}(d))$ times. ∎